# Natural Language Actor-Critic: Scalable Off-Policy Learning in Language Space

### v1

@@ Abstract
Large language model (LLM) agents—LLMs that dynamically interact with an environment over long horizons—have become an increasingly important area of research, enabling automation in complex tasks involving tool-use, web browsing, and dialogue with people. In the absence of expert demonstrations, training LLM agents has relied on policy gradient methods that optimize LLM policies with respect to an (often sparse) reward function. However, in long-horizon tasks with sparse rewards, learning from trajectory-level rewards can be noisy, leading to training that is unstable and has high sample complexity. Furthermore, policy improvement hinges on discovering better actions through exploration, which can be difficult when actions lie in natural language space. In this paper, we propose Natural Language Actor-Critic (NLAC), a novel actor-critic algorithm that trains LLM policies using a generative LLM critic that produces natural language rather than scalar values. This approach leverages the inherent strengths of LLMs to provide a richer and more actionable training signal; particularly, in tasks with large, open-ended action spaces, natural language explanations for why an action is suboptimal can be immensely useful for LLM policies to reason how to improve their actions, without relying on random exploration. Furthermore, our approach can be trained off-policy without policy gradients, offering a more data-efficient and stable alternative to existing on-policy methods. We present results on a mixture of reasoning, web browsing, and tool-use with dialogue tasks, demonstrating that NLAC shows promise in outperforming existing training approaches and offers a more scalable and stable training paradigm for LLM agents.
@@ 1 Introduction
While LLMs excel at natural language tasks like question-answering (pyatkin2022reinforced) and problem-solving (hendrycksmath2021; jimenez2024swebench), which can be solved with a single response, LLM agent tasks require multi-turn interactions. Specifically, LLM agent tasks require the model to act within an environment, by taking actions sequentially and observing their results, ultimately to accomplish some long-term goal. Such tasks include autonomous reasoning (deepresearch), tool-use (nakano2022webgpt), and dialogue with users (hong2023zeroshot; yu2023promptbasedmontecarlotreesearch). These tasks require agents to dynamically plan and intelligently respond to environmental stimuli, which base, pretrained LLMs struggle to do without additional training (bachmann2024pitfalls). To train effective LLM agents, we will need algorithms that can fine-tune LLMs to pursue temporally extended goals in the context of multi-turn, long-horizon interactions.
Currently, LLM agents are trained with a variety of methods, often combining supervised fine-tuning (SFT) with reinforcement learning (RL) (rafailov2023direct; carta2023grounding; wang2025stecastepleveltrajectorycalibration). For complex agentic tasks where labeled expert data is expensive to collect, such as ones involving interaction with real users, the prevailing training methods focus on policy optimization using algorithms such as Proximal Policy Optimization (PPO)  (schulman2017proximal) or Group Relative Policy Optimization (GRPO) (shao2024deepseekmathpushinglimitsmathematical). The LLM agents are trained to generate environment actions accompanied by high-level reasoning to explain their decision-making (yao2022react; wei2023chainofthought). These methods are designed to teach an LLM to reason about the problem, plan over appropriate actions, and learn from the environment observations. However, prior attempts of using RL to train LLM agents present significant problems. First, these algorithms are notoriously data-inefficient because they are on-policy, meaning they require sampling new trajectories from the current policy at every training step. Second, and perhaps more importantly, they rely on an often sparse, scalar reward as their only training signal, which can be a weak and unstable signal for learning robust, generalizable strategies over long-horizon tasks.
In this work, we propose a new actor-critic algorithm (haarnoja2018soft) to train LLM agents, where a critic (which estimates the value of actions) is jointly learned with a policy, both using off-policy data. In contrast to traditional actor-critic to train LLM agents (qtransformer; zhou2024archertraininglanguagemodel), we believe training a critic that outputs textual evaluations more effectively leverages the text-based reasoning capabilities of pretrained LLMs. Specifically, policy optimization using scalar values requires the policy to discover actions of high value through random exploration. If values were instead in natural language space, an LLM policy could understand how to improve its decision-making, reducing the reliance on random chance to uncover better actions.
Prior methods exist that train critics to generate language evaluations (feng2025naturallanguagereinforcementlearning; hong2025planningsearchrefiningfrontier). Notably, feng2025naturallanguagereinforcementlearning propose Natural Language Reinforcement Learning (NLRL) as a framework for learning policies and critics in language space. Our work aims to address key limitations in NLRL to make policy iteration in language space scalable to all LLM agent tasks. Specifically, NLRL relies on enumerating over both environment transitions, as well as potential actions, and aggregating them using in-context learning. We believe such training is impractical for tasks with complex dynamics or action spaces (such as dialogue), as enumerating over all possibilities and fitting them in-context is infeasible with limited time and memory. Our algorithm falls under the paradigm introduced by NLRL, but proposes novel objectives to train the critic and improve the policy that scales to learning general LLM agents.
@@
In this paper, we propose Natural Language Actor-Critic (NLAC), a novel algorithm for training LLM agents, where a natural language critic is jointly trained with a policy, and its evaluations directly inform how to perform policy improvement. Theoretically, we are able to connect the learned representations of our critic to successor features (barreto2018successorfeaturestransferreinforcement), allowing us to prove convergence to the optimal policy. Empirically, we evaluate our approach on a range of LLM agent tasks, ranging from reasoning, tool-use, and dialogue. Our empirical results demonstrate substantial improvement over prior approaches to learn LLM agents, showing our algorithm is an appealing alternative to prevailing on-policy training methods.
@@ 2 Related Work
LLM agents. LLM agents can be used to tackle a variety of complex real-world tasks, including dialogue (hong2023zeroshot; yu2023promptbasedmontecarlotreesearch), tool-use (nakano2022webgpt; schick2023toolformerlanguagemodelsteach), and embodied decision-making  (wang2023voyager). The primary challenge in the design of effective LLM agents is enabling LLMs, which traditionally excel at generating single-step responses, to interact sequentially with an environment to accomplish a long-term objective. ReAct prompting is a popular method to leverage chain-of-thought reasoning of LLMs for long-horizon planning, by instructing LLMs to explicitly articulate their high-level plans (yao2022react). More recent approaches have explored the capability of LLM agents to self-correct their initial attempts at planning using more sophisticated prompting techniques  (shinn2023reflexionlanguageagentsverbal; madaan2023selfrefineiterativerefinementselffeedback; zhou2024languageagenttreesearch). For example, Reflexion prompting adds a step of self-reflection on top of ReAct to allow LLM agents to refine their initial reasoning after some environment feedback  (shinn2023reflexionlanguageagentsverbal). However, self-correction methods rely the ability to “backtrack," or undo previous actions, whereas we measure the capability of LLM agents with one chance to solve a task.
@@
Process reward models. One of the primary challenges in learning LLM agents is the reliance on a single, sparse reward for long-horizon interactions. This makes credit assignment, or distinguishing between good and bad actions in a long rollout, difficult. Process reward models (PRMs) aim to address this, particularly by providing action-level feedback using either human annotations  (lightman2023letsverifystepstep), or an estimated value function in the absence of human intervention (wang2024mathshepherdverifyreinforcellms; setlur2024rewardingprogressscalingautomated). Our learned natural language critic can be considered an instance of an PRM, but unlike traditional PRMs that provide scalar feedback over actions, our critic outputs feedback in language space. We believe such feedback is more useful for LLM policies that can understand and articulate their decisions in natural language.
@@
Reinforcement learning for LLM agents. More recently, multiple works have attempted to explicitly fine-tune LLMs as agents using RL (carta2024groundinglargelanguagemodels; zhou2024archertraininglanguagemodel). The primary way this was done was naively adapting traditional RL fine-tuning used to align LLM responses to multi-turn tasks with environment interaction (stiennon2020learning; ouyang2022training; ramamurthy2023is). These methods used PPO (schulman2017proximal) to finetune LLMs using the environment reward. However, traditional policy optimization for long-horizon tasks exacerbates the instabilities of RL training, particularly due to reliance on exploration and proper credit assignment. In this work, we hypothesize that training in natural language over scalar space improves stability and sample efficiency, particularly in better leveraging the capabilities of LLMs to understand and articulate thoughts in natural language. The closest work to ours that does this is NLRL (feng2025naturallanguagereinforcementlearning), which also proposes learning value functions that output text. However, in NLRL, these values are obtained via repeated sampling of on-policy trajectories and aggregating them in-context. In addition, policy improvement is achieved by enumerating over possible actions and their evaluations. We believe such enumeration and aggregation in-context is intractable for tasks with complex dynamics and large action spaces. Our method circumvents these drawbacks by training the critic to probabilistically generate textual rollouts via a novel language Bellman backup, and treating policy improvement as iterative refinement.
@@ 3 Preliminaries
Markov decision processes. We adopt the formalism of a Markov decision process (MDP) given by $M=({\mathcal{S}},{\mathcal{A}},P,r,\rho,\gamma)$, where $\mathcal{S}$} is the state space, $\mathcal{A}$} is the action space, $P$ is the transition function, $r$ is the reward function, $\rho$ is the initial state distribution, and $\gamma$ is the discount factor. When action $a\in{\mathcal{A}}$ is executed at state $s\in{\mathcal{S}}$, the next state is sampled $s^{\prime}\sim P(\cdot|s,a)$, and the agent receives reward $r$ with mean $r(s,a)\in\mathbb{R}$.
@@
ReAct prompting. LLM agents are commonly implemented using ReAct prompting to better leverage the base reasoning capabilities of LLMs  (yao2022react). ReAct prompting instructs LLM agents to output actions $a_{t}\sim\pi(\cdot|s_{t})$ that are actually composite, consisting of a thought $a_{t}^{\mathsf{tht}}$ where the agent performs a reasoning step, followed by the actual environment action $a_{t}^{\mathsf{env}}$. For example, in dialogue, the thought could be the high-level strategy or plan the agent aims to execute, whereas the environment action is the actual utterance by the agent. Then, the transition function appends to $s_{t}$ the environment action $a_{t}^{\mathsf{env}}$ as well as any new observations by the environment $o_{t+1}$, to form the next state $s_{t+1}$. Note that the thought does not affect the transition dynamics, namely $P(\cdot|s_{t},a_{t})=P(\cdot|s_{t},a_{t}^{\mathsf{env}})$.
Reinforcement learning. The objective of RL is to find a policy $\pi$ that maximizes the expected discounted return $J(\pi)=\mathbb{E}_{\tau\sim p^{\pi}}\left[\sum_{t=0}^{T-1}\gamma^{t}r(s_{t},a_{t})\right]$ in an MDP, where $\tau=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{T})$ and $p^{\pi}(\tau)=\rho(s_{0})\prod_{t=0}^{T-1}\pi(a_{t}|s_{t})P(s_{t+1}|s_{t},a_{t})$. Standard policy gradient approaches directly train policy $\pi$ using the gradient of $\nabla_{\pi}J(\pi)$, while more sophisticated algorithms such as PPO and GRPO additionally clip the updates to improve stability  (schulman2017proximal; shao2024deepseekmathpushinglimitsmathematical). Actor-critic algorithms additionally learn a state-action value function, or Q-function, defined as $Q^{\pi}(s_{t},a_{t})=\mathbb{E}_{(s,a)_{t+1:\infty}\sim p^{\pi}}\left[\sum_{t^{\prime}=t}^{T-1}\gamma^{t^{\prime}-t}r(s_{t^{\prime}},a_{t^{\prime}})\right]$. Such Q-functions are learned by regressing to their Bellman backup:
$\mathcal{B}Q^{\pi}(s_{t},a_{t})=r(s_{t},a_{t})+\mathbb{E}_{s_{t+1},a_{t+1}\sim P^{\pi}}\left[Q^{\pi}(s_{t+1},a_{t+1})\right]\,$,
where $P^{\pi}(s^{\prime},a^{\prime}|s,a)=P(s^{\prime}|s,a)\pi(a^{\prime}|s^{\prime})$. Then, an improved policy $\pi^{\prime}$ can be derived using the Q-function via greedy or maximum-entropy optimization $\pi^{\prime}(a_{t}|s_{t})\propto\exp(Q^{\pi}(s_{t},a_{t}))$.
@@ 4.1 Policy Evaluation
Predicting the future using language. The key addition that is not captured by scalar Q-values is an explanation of why a particular action is optimal or not. As we will discuss later, this information is valuable for LLM policies to understand how to refine their actions during policy improvement, avoiding the reliance on random exploration to discover better actions. We believe that the key for a critic to derive these explanations is the prediction and analysis of future outcomes. In order to do so, we must train our natural language critic to additionally act as a successor function, defined as follows:
@@
A language successor model $M^{\pi}$ for policy $\pi$ takes a state $s_{t}$ and action $a_{t}$ as input, and probabilistically generates a textual description of rollout $(s,a)_{t+1:T}$, or what will happen to policy $\pi$ in the future, and reward $r(s_{T})$. We denote by $M^{\pi}(\cdot\mid s_{t},a_{t})$ the distribution from which such descriptions are sampled.
Our language successor model shares similarities with successor features (barreto2018successorfeaturestransferreinforcement) in that both can predict a distribution over future rollouts, and—as we show later—be trained using temporal difference learning. The main difference lies in that traditional successor features are used to compute Q-values via a linear product, whereas ours is used to generate state-action values in natural language via output by an LLM.
To train our language successor model, we draw inspiration from distributional value learning (bellemare2017distributional), which introduces a distributional Bellman backup to train a distribution over returns rather than just their scalar expectation. Notably, the distributional Bellman backup used one-step samples of the future and thus could be computed off-policy. Similarly, we propose a language Bellman backup $\mathcal{B}_{L}$ that bears some semblance to the distributional Bellman backup, but makes key adaptations to account for samples that are textual descriptions of rollouts rather than scalar returns.
@@
$\qquad\qquad\mathcal{B}_{L}\,M^{\pi}(d_{t}\mid s_{t},a_{t})$ $=\mathsf{Pr}\left[d_{t}=B(r(s_{t},a_{t}),s_{t+1},a_{t+1},d_{t+1})\mid s_{t+1},a_{t+1},d_{t+1}\right]\,$, (1)
$\quad s_{t+1},a_{t+1}\sim P^{\pi}(\cdot\mid s_{t},a_{t})\,,\ d_{t+1}\sim M^{\pi}(\cdot\mid s_{t+1},a_{t+1})\,$,
where $B$ is a function that combines immediate next state and action $s_{t+1},a_{t+1}$ with description $d_{t+1}$ of rollout $(s,a)_{t+2:T}$ into one description of the concatenated rollout $(s,a)_{t+1:T}$.
@@
Then, we can train our language successor model $M^{\pi}$ by minimizing the divergence between distributions $M^{\pi}(\cdot|s_{t},a_{t})$ and target distributions created by the language Bellman backup:
$M^{\pi}$ $=\arg\min_{M}\mathbb{E}_{(s_{t},a_{t}s_{t+1})\sim\mathcal{D}}\left[D_{f}\left(M(\cdot\mid s_{t},a_{t})\ ||\ \mathcal{B}_{L}\,M(\cdot\mid s_{t},a_{t})\right)\right]\,.$ (2)
@@
Then, we can approximate natural language value $Q^{\pi}_{L}(s_{t},a_{t})$ as:
$\qquad\qquad\qquad Q^{\pi}_{L}(s_{t},a_{t})$ $\approx E(s_{t},a_{t},d_{t}^{(1)},\ldots,d_{t}^{(k)})\,,\quad d_{t}^{(i)}\sim M^{\pi}(\cdot\mid s_{t},a_{t})\,,\ \forall\,i\in[k]\,.$ (3)
Note that $E$ essentially aggregates and summarizes multiple descriptions of different rollouts that are all fit in-context, which LLMs have demonstrated a priori efficacy in without additional training (feng2025naturallanguagereinforcementlearning). This means that the only training required to perform evaluation of policy $\pi$ in language space is learning the language successor model $M^{\pi}$ (see Figure 1 for illustration).

### v2

@@ Abstract
Large language model (LLM) agents—LLMs that dynamically interact with an environment over long horizons—have become an increasingly important area of research, enabling automation in complex tasks involving tool-use, web browsing, and dialogue with people. In these settings, traditional policy gradient methods may suffer from unstable learning and poor sample-complexity due to poor credit assignment. Meanwhile, actor-critic methods address long horizons by learning a critic that provides more granular feedback and enables off-policy (and potentially offline) learning, but are heavily dependent on the fidelity of the critic. In this paper, we propose Natural Language Actor-Critic (NLAC), a novel actor-critic algorithm that trains LLM policies using a generative LLM critic that produces values in natural language space. While natural language values have been leveraged in the past to provide a more flexible and actionable training signal, our work is the first to perform general and scalable value learning without relying on in-context aggregation of information from on-policy rollouts. This means our approach can be trained off-policy without policy gradients, offering a more sample-efficient alternative to existing methods. We present results on a mixture of reasoning, web browsing, and tool-use with dialogue tasks, demonstrating that NLAC shows promise in outperforming existing training approaches and offers a more scalable and stable training paradigm for LLM agents.
Machine Learning, ICML
@@ 1 Introduction
While LLMs excel at natural language tasks like question-answering (Pyatkin et al., 2022) and problem-solving (Hendrycks et al., 2021; Jimenez et al., 2024), which can be solved with a single response, LLM agent tasks require multi-turn interactions. Such tasks include autonomous reasoning (OpenAI, 2025), tool-use (Nakano et al., 2022), and dialogue with users (Hong et al., 2023; Yu et al., 2023). These tasks require agents to dynamically plan and respond to environmental stimuli, which base LLMs struggle to do without additional fine-tuning (Bachmann & Nagarajan, 2024). To train effective LLM agents, we will need algorithms that can fine-tune LLMs to pursue temporally extended goals within multi-turn interactions. However, in long-horizon settings where sparse rewards offer poor credit assignment, training LLM policies using standard policy gradient approaches can be unstable and inefficient.
Actor-critic methods tackle long-horizon tasks by learning a critic that evaluates actions and provides a denser training signal, improving sample efficiency and enabling off-policy or offline learning. However, such methods are highly dependent on the quality of the critic, and prior attempts at actor-critic methods for LLM policies have not demonstrated convincing improvement over standard fine-tuning techniques (Sodhi et al., 2023; Abdulhai et al., 2023). We posit that this is because standard scalar-valued critics are a poor fit for language action spaces: they compress rich failure modes into a single scalar and provide little guidance on how a language policy should improve its actions.
To address the drawbacks of standard critics for language policies, Feng et al. (2025) propose Natural Language Reinforcement Learning (NLRL), which frames policy evaluation and improvement directly in language space. Natural language values can provide richer, more flexible credit assignment than scalar critics, since they can explain why an action is good or bad and suggest how it should be refined. However, existing approaches to NLRL typically rely on aggregating information from many on-policy rollouts in-context. This becomes increasingly unscalable as trajectories grow longer and environment dynamics become more complex; hence, thus far NLRL has largely been evaluated in simple environments such as gridworld navigation.
@@
In this paper, we propose Natural Language Actor-Critic (NLAC), a novel algorithm for training LLM agents, where a natural language critic is jointly trained with a policy in a general and scalable manner. The key insight that allows NLAC to address the limitations of NLRL is the training of the critic to predict infinite-horizon futures as compact summaries. This is done via a novel language Bellman backup objective that bootstraps from one-step transitions, and completely circumvents in-context aggregation over on-policy samples and enables scalable off-policy value learning in language space. Theoretically, we are able to connect the learned representations of our critic to successor features (Barreto et al., 2017), allowing us to prove convergence to the optimal policy. Empirically, we evaluate our approach on a range of LLM agent tasks, ranging from reasoning, tool-use, and dialogue. Our empirical results demonstrate substantial improvement over prior approaches to learn LLM agents, showing our algorithm is an appealing alternative to prevailing on-policy training methods.
@@ 2 Related Work
LLM agents. LLM agents can be used to tackle a variety of complex real-world tasks, including dialogue (Hong et al., 2023; Yu et al., 2023), tool-use (Nakano et al., 2022; Schick et al., 2023), and embodied decision-making  (Wang et al., 2023). The primary challenge in the design of effective LLM agents is enabling LLMs, which traditionally excel at generating single-step responses, to interact sequentially with an environment to accomplish a long-term objective. ReAct prompting is a popular method to leverage chain-of-thought reasoning of LLMs for long-horizon planning, by instructing LLMs to explicitly articulate their high-level plans (Yao et al., 2022). More recent approaches have explored the capability of LLM agents to self-correct their initial attempts at planning using more sophisticated prompting techniques  (Shinn et al., 2023; Madaan et al., 2023; Zhou et al., 2024a). For example, Reflexion prompting adds a step of self-reflection on top of ReAct to allow LLM agents to refine their initial reasoning after some environment feedback  (Shinn et al., 2023). However, self-correction methods rely the ability to “backtrack," or undo previous actions, whereas we measure the capability of LLM agents with one chance to solve a task.
@@
Process reward models. One of the primary challenges in learning LLM agents is the reliance on a single, sparse reward for long-horizon interactions. This makes credit assignment, or distinguishing between good and bad actions in a long rollout, difficult. Process reward models (PRMs) aim to address this, particularly by providing action-level feedback using either human annotations  (Lightman et al., 2023), or an estimated value function in the absence of human intervention (Wang et al., 2024; Setlur et al., 2025). Our learned natural language critic can be considered an instance of an PRM, but unlike traditional PRMs that provide scalar feedback over actions, our critic outputs feedback in language space. We believe such feedback is more useful for LLM policies that can understand and articulate their decisions in natural language.
@@
Reinforcement learning for LLM agents. More recently, multiple works have attempted to explicitly fine-tune LLMs as agents using RL (Carta et al., 2024; Zhou et al., 2024b). The primary way this was done was naively adapting traditional RL fine-tuning used to align LLM responses to multi-turn tasks with environment interaction (Stiennon et al., 2020; Ouyang et al., 2022; Ramamurthy et al., 2023). These methods used PPO (Schulman et al., 2017) to finetune LLMs using the environment reward. However, traditional policy optimization for long-horizon tasks exacerbates the instabilities of RL training, particularly due to reliance on exploration and proper credit assignment. In this work, we hypothesize that training in natural language over scalar space improves stability and sample efficiency, particularly in better leveraging the capabilities of LLMs to understand and articulate thoughts in natural language. The closest work to ours that does this is NLRL (Feng et al., 2025), which also proposes learning value functions that output text. However, in NLRL, these values are obtained via repeated sampling of on-policy trajectories and aggregating them in-context. In addition, policy improvement is achieved by enumerating over possible actions and their evaluations. We believe such enumeration and aggregation in-context is intractable for tasks with complex dynamics and large action spaces. Our method circumvents these drawbacks by training the critic to probabilistically generate compact summaries of rollouts via a novel language Bellman backup, then updating a policy through distillation from a refinement policy that conditions on the critic output.
@@ 3 Preliminaries
Markov decision processes. We adopt the formalism of a Markov decision process (MDP) given by $M=(\mathcal{S},\mathcal{A},P,r,\rho,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P$ is the transition function, $r$ is the reward function, $\rho$ is the initial state distribution, and $\gamma$ is the discount factor. When action $a\in\mathcal{A}$ is executed at state $s\in\mathcal{S}$, the next state is sampled $s^{\prime}\sim P(\cdot|s,a)$, and the agent receives reward $r(s,a)\in\mathbb{R}$.
@@
ReAct prompting. LLM agents are commonly implemented using ReAct prompting to better leverage the base reasoning capabilities of LLMs  (Yao et al., 2022). ReAct prompting instructs LLM agents to output actions $a_{t}\sim\pi(\cdot|s_{t})$ that are actually composite, consisting of a thought $a_{t}^{\mathsf{tht}}$ where the agent performs a reasoning step, followed by the actual environment action $a_{t}^{\mathsf{env}}$. For example, in dialogue, the thought could be the high-level strategy or plan the agent aims to execute, whereas the environment action is the actual utterance by the agent. Then, the transition function appends to $s_{t}$ the environment action $a_{t}^{\mathsf{env}}$ as well as any new observations by the environment $o_{t+1}$, to form the next state $s_{t+1}$. Note that the thought does not affect the transition dynamics, namely $P(\cdot|s_{t},a_{t})=P(\cdot|s_{t},a_{t}^{\mathsf{env}})$.
Reinforcement learning. The objective of RL is to find a policy $\pi$ that maximizes the expected discounted return $J(\pi)=\mathbb{E}_{\tau\sim p^{\pi}}\left[\sum_{t=0}^{T-1}\gamma^{t}r(s_{t},a_{t})\right]$ in an MDP, where $\tau=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{T})$ and $p^{\pi}(\tau)=\rho(s_{0})\prod_{t=0}^{T-1}\pi(a_{t}|s_{t})P(s_{t+1}|s_{t},a_{t})$. Standard policy gradient approaches directly train policy $\pi$ using the gradient of $\nabla_{\pi}J(\pi)$, while more sophisticated algorithms such as PPO and GRPO additionally clip the updates to improve stability  (Schulman et al., 2017; Shao et al., 2024). Actor-critic algorithms additionally learn a state-action value function, or Q-function, defined as $Q^{\pi}(s_{t},a_{t})=\mathbb{E}_{(s,a)_{t+1:\infty}\sim p^{\pi}}\left[\sum_{t^{\prime}=t}^{T-1}\gamma^{t^{\prime}-t}r(s_{t^{\prime}},a_{t^{\prime}})\right]$. Such Q-functions are learned by regressing to their Bellman backup:
$\mathcal{B}Q^{\pi}(s_{t},a_{t})$ $=r(s_{t},a_{t})+\mathbb{E}_{s_{t+1},a_{t+1}\sim P^{\pi}}\!\left[Q^{\pi}(s_{t+1},a_{t+1})\right].$
where $P^{\pi}(s^{\prime},a^{\prime}|s,a)=P(s^{\prime}|s,a)\pi(a^{\prime}|s^{\prime})$. Then, an improved policy $\pi^{\prime}$ can be derived using the Q-function via greedy or maximum-entropy optimization.
@@ 4.1 Policy Evaluation
Predicting the future using language. The key addition that is not captured by scalar Q-values is an explanation of why a particular action is optimal or not. As we will discuss later, this information is valuable for LLM policies to understand how to refine their actions during policy improvement, avoiding the reliance on random exploration to discover better actions. We believe that the key for a critic to derive these explanations is the prediction and analysis of future outcomes. However, full trajectories can be prohibitively complex to model, so we instead predict compact summaries of futures. In order to do so, we must train our natural language critic to additionally act as a successor function, defined as follows:
@@
A language successor model $M^{\pi}$ for policy $\pi$ takes a state $s_{t}$ and action $a_{t}$ as input, and probabilistically generates a textual summary of rollout $(s,a)_{t+1:T}$, or what will happen to policy $\pi$ in the future, and reward $r(s_{T})$. We denote by $M^{\pi}(\cdot\mid s_{t},a_{t})$ the distribution from which such descriptions are sampled.
Our language successor model shares similarities with successor features (Barreto et al., 2017) in that both can predict a distribution over future rollouts, and—as we show later—be trained using temporal difference learning. The main difference lies in that traditional successor features are used to compute Q-values via a linear product, whereas ours is used to generate state-action values in natural language via output by an LLM.
To train our language successor model, we draw inspiration from distributional value learning (Bellemare et al., 2017), which introduces a distributional Bellman backup to train a distribution over returns rather than just their scalar expectation. Notably, the distributional Bellman backup used one-step samples of the future and thus could be computed off-policy. Similarly, we propose a language Bellman backup $\mathcal{B}_{L}$ that bears some semblance to the distributional Bellman backup, but makes key adaptations to account for samples that are textual descriptions of rollouts rather than scalar returns.
@@
$\mathcal{B}_{L}\,M^{\pi}(d_{t}\mid s_{t},a_{t})=$ (1)
$\mathsf{Pr}\Bigl[d_{t}=B\bigl(r(s_{t},a_{t}),s_{t+1},a_{t+1},d_{t+1}\bigr)\ \Big|\ \ s_{t+1},a_{t+1},d_{t+1}\Bigr]\,$,
$\quad s_{t+1},a_{t+1}\sim P^{\pi}(\cdot\mid s_{t},a_{t})\,,\ d_{t+1}\sim M^{\pi}(\cdot\mid s_{t+1},a_{t+1})\,.$
where $B$ is a function that combines immediate next state and action $s_{t+1},a_{t+1}$ with description $d_{t+1}$ of rollout $(s,a)_{t+2:T}$ into one concatenated description of $(s,a)_{t+1:T}$.
@@
Then, we can train our language successor model $M^{\pi}$ by minimizing the divergence between distributions $M^{\pi}(\cdot|s_{t},a_{t})$ and target distributions created by the language Bellman backup: $M^{\pi}=$
$M^{\pi}=$ (2)
$\arg\min_{M}\mathbb{E}_{(s_{t},a_{t},s_{t+1})\sim\mathcal{D}}\Bigl[D_{f}\bigl(M(\cdot|s_{t},a_{t})\|\mathcal{B}_{L}M(\cdot|s_{t},a_{t})\bigr)\Bigr]\,.$
@@
Then, we can approximate $Q^{\pi}_{L}(s_{t},a_{t})$ as:
$Q^{\pi}_{L}(s_{t},a_{t})\approx E(s_{t},a_{t},d_{t}^{(1)},\ldots,d_{t}^{(k)})\,$, (3)
$\quad d_{t}^{(i)}\sim M^{\pi}(\cdot\mid s_{t},a_{t})\,,\ \forall\,i\in[k]\,.$
Note that $E$ essentially aggregates and summarizes multiple descriptions of different rollouts that are all fit in-context, which LLMs have demonstrated a priori efficacy in without additional training (Feng et al., 2025). This means that the only training required to perform evaluation of policy $\pi$ in language space is learning the language successor model $M^{\pi}$ (see Figure 1 for illustration).